Adult tickets to a show are $9. Student tickets are $6. If you sold 29 tickets for a total of $210, how many adult tickets did you sell
only use 1 variable

Answer by MathPro(14)

Good Afternoon,
Your problem is:
Adult tickets to a show are $9. Student tickets are $6. If you sold 29 tickets for a total of $210, how many adult tickets did you sell
only use 1 variable
We are given that the total amount of tickets sold is 29. In words this means that the number of adult tickets + the number of student tickets = 29
I’ll digress for a moment and ask you to imagine if the total number of tickets was 3 and we were told that the number of adult tickets sold was 1 , then that would mean the number of student tickets sold would be 3 – 1 = 2.
Let’s apply this to our problem. Call the number of adult tickets sold, x, then the number of student tickets sold is 29 – x.
I’ll just digress and say that if you bought two items at 50 cents each the total cost would be 2($0.50) = $1.00
We are told that adult tickets are $9.00 each, so the total cost of all the adult tickets together is $9(x) and similarly the total cost of the student tickets is $6(29 – x). We are given that the total price of the adult and student tickets combined is $210.00. We can write this as:
$9x + $6(29 – x) = $210 equation (1)
Multiply out the second term:
$9x + $6(29) - $6x = $210
$3x + $174 = $210
Subtract $174 from both sides
$3x = $210 - $174 = $36
Divide both sides by $3:
$3x/$3 = $36/$3
In the same way that 3/3 = 1, $/$ = 1. So we get
x = 12
This means 12 adult tickets was sold and (29 -12) = 17 student tickets was sold
Let’s check the answer by substituting for x into equation (1)
$9(12) + $6(17) = $210
The left side is equal to the right side so our answer is correct.
Good Luck!
Selma from Rent-A-Whiz, an online tutoring and homework help website. Just visit www.rentawhiz.com to learn more. Currently, we are offering to solve one math problem for free. Just visit our website, select the homework tab, then click on the math problem submission link and quote coupon code 003 when you submit your problem.